Problem 5, 14 pts] With the TFSA limit at $6,000 for this year, the total room available in 2019 for you since you have never contributed and have been eligible for the TFSA since its introduction in 2009 is $63,500. Thus, you decide to put 20% of your money into a stock with an expected return of 12%, $30,500 into a stock with an expected return of 25%, and the rest into a stock with an expected return of 20%. What is the expected return of your portfolio? (Show, your calculations)
Answer #1

Answer:

Total Investable amount I=$63500

Investment in stock 1 w1=20%

Return from stock 1 r1=12%

Investment in stock 2 w2=30500/63500=48.0%

Return from stock 2 r2=25%

Investment in stock 3 w3=100%-20%-48.0%=32%

Return from stock 3 r3=20%

Expected return from portfolio = w1*r1+w2*r2+w3*r3=20%*12%+48%*25%+32%*20%=20.80%